2) To start an avalanche on a mountain slope, an artillery shell is fired with an initial velocity of 300 m/s at 55.0° above the horizontal. It explodes on the mountainside 42.0 s after firing. What are the x and y coordinates of the shell where it explodes, relative to its firing point?
1
Expert's answer
2019-11-24T17:10:35-0500
"x=v\cos{55}t=300(42)\cos{55}=7230\ m"

"y=v\sin{55}t-0.5gt^2=300(42)\sin{55}-0.5(9.8)(42)^2=1680\ m"